The 3th overtone of one end fixed string produce a frequency of 72Hz and the
length between a node and the adjacent anti-node is 5cm.waht is:
(a) the length of the string ?
1
Expert's answer
2019-11-05T10:48:37-0500
"\frac{\lambda_3}{4}=5\ cm\to \lambda_3=20\ cm"

"L=\frac{3}{2}\lambda_3=\frac{3}{2}20=30\ cm."